OpenStudy (anonymous):

Which of the binomials below is a factor of this trinomial? x2 - 9x + 18 A. x - 6 B. x + 3 C. x + 6 D. x - 9

OpenStudy (amoodarya):

turn into form of (x-a)^2+k

OpenStudy (amoodarya):

\[x^2 - 9x + 18=(x-\frac{ 9 }{ 2 })^2-(\frac{ 9 }{ 2 })^2+18 =\\\]

OpenStudy (anonymous):

so whats the answer? A?B?C?D? Thanks

OpenStudy (anonymous):

because i dont really get this topic

OpenStudy (amoodarya):

another way is (x-...)(x-...)=x^2-9x+18 (x-3)(x-6)

OpenStudy (amoodarya):

-3-6=-9 -3*(-6)=18

OpenStudy (anonymous):

How about this one. Which of the binomials below is a factor of this trinomial? x2 - x - 12 A. x - 6 B. x - 3 C. x + 6 D. x + 3